Check The following is the balance sheet and income statement for Metro Eagle Outfitters, in condensed form, plus some informRequired: Calculate the following liquidity ratios for Metro Eagle in 2018 and 2019: 1. Inventory turnover 2. Current ratio.
Answer #1

2019

1) inventory turnover = cost of goods sold / average inventory

= 2,102,480 / 375,483

= 5.60 times

2) current ratio = current assets / current liabilities

= 1,158,800 / 475,402

= 2.44

3) quick ratio = current assets - inventory / current liabilities

= 1,158,800 - 349,452 / 475,402

= 1.70

4) cash flow ratio = cash flow from operation / current liabilities

= 593,171 / 475,402

= 1.25

2018

1) inventory turnover = cost of goods sold / average inventory

= 1,992,471 / 359,861

= 5.54 times

2) current ratio = current assets / current liabilities

= 1,321,488 / 456,401

= 2.90

3) quick ratio = current assets - inventory / current liabilities

= 1,321,488 - 401,514 / 456,401

= 2.01

4) cash flow ratio = cash flow from operation / current liabilities

= 508,817 / 456,401

= 1.11